# Gödel Agent: A Self-Referential Agent Framework for Recursively Self-Improvement

### v3

@@ Abstract
The rapid advancement of large language models (LLMs) has significantly enhanced the capabilities of agents across various tasks. However, existing agentic systems, whether based on fixed pipeline algorithms or pre-defined meta-learning frameworks, cannot search the whole agent design space due to the restriction of human-designed components, and thus might miss the more optimal agent design. In this paper, we introduce Gödel Agent, a self-evolving framework inspired by the Gödel machine, enabling agents to recursively improve themselves without relying on predefined routines or fixed optimization algorithms. Gödel Agent leverages LLMs to dynamically modify its own logic and behavior, guided solely by high-level objectives through prompting. Experimental results on multiple domains demonstrate that implementation of Gödel Agent can achieve continuous self-improvement, surpassing manually crafted agents in performance, efficiency, and generalizability. Our code can be found at https://github.com/Arvid-pku/Godel_Agent.
@@
Xunjian Yin111 Work completed during internship at UCSB. , Xinyi Wang , Liangming Pan , Li Lin Xiaojun Wan , William Yang Wang Peking University  University of California, Santa Barbara   University of Arizona {xjyin,efsotr_l,wanxiaojun}@pku.edu.cnxinyi_wang@ucsb.edu liangmingpan@arizona.eduwilliam@cs.ucsb.edu
@@ 1 Introduction
There is a large body of work proposing agents capable of self-refinement. Some agents are designed to iterate over a fixed routine consisting of a list of fixed modules, while some of the modules are capable of taking self- or environment feedback to refine their actions  (Chen et al., 2023b; Qu et al., 2024a; Yao et al., 2023). This type of agent, referred to as Hand-Designed Agent, is depicted as having the lowest degree of freedom in Figure 2. More automated agents have been designed to be able to update their routines or modules in some pre-defined meta-learning routine, for example, natural language gradients (Zhou et al., 2024), meta agent (Hu et al., 2024), or creating and collecting demonstrations  (Khattab et al., 2023). This type of agent, known as Meta-Learning Optimized Agents, is depicted as having the middle degree of freedom in Figure 2. However, there are inevitably some human priors involved in these agent designs that cannot be improved during the inference time.
@@
In this paper, we choose to implement it by letting it manipulate its own runtime memory, i.e., the agent is able to retrieve its current code in the runtime memory and modify it by monkey patching (Bimal, 2012), which dynamically modifies classes or modules during execution. To allow it to update the logic of the running main function, unlike the loop-iterative approach of traditional agents, we implement the main function as a recursive function. In this function, LLM analyzes and makes a series of decisions, including reading and modifying its own code from runtime memory (self-awareness***In this paper, self-awareness means that the agent can introspect and read its own code and files, not to imply any philosophical sense of consciousness or awareness. and self-modification), and interacting with the environment to gather feedback. The agent then proceeds to the subsequent recursive depth and continues to optimize itself.
To validate the effectiveness of Gödel Agent, we conduct experiments on multiple domains including coding, science, math, and reasoning. Our results demonstrate that Gödel Agent achieves significant performance gain across various tasks, surpassing various widely-used agents that require human design. The same implementation of Gödel Agent can easily adapt to different tasks by only specifying the environment description and feedback mechanism. Additionally, the case study of the optimization progress reveals that Gödel Agent can provide novel insights into agent design.
@@ 4.2 Additional Designs
Table 1: Results of three paradigms of agents on different tasks. The highest value is highlighted in bold, and the second-highest value is underlined. Gödel-base is the constrained version of Gödel Agent, allowing for fair comparisons with other baselines. Gödel-free represents the standard implementation without any constraints, whose results are italicized. We report the test accuracy and the 95% bootstrap confidence interval on test setsThe results of baseline models are refer to Hu et al. (2024)..
@@ 6.3 Case Study: Game of 24
The findings indicate that stronger initial policies lead to faster convergence, with smaller optimization margins, as Gödel Agent reaches its performance limit without further enhancing its optimization capabilities. Conversely, weaker initial methods result in slower convergence and larger gains, with Gödel Agent making more modifications. However, even in these cases, Gödel Agent does not outperform the results achieved using ToT. Given the current limitations of LLMs, it is challenging for Gödel Agent to innovate beyond state-of-the-art algorithms. Improvements in LLM capabilities are anticipated to unlock more innovative self-optimization strategies in the future. We also discuss the future directions in Appendix E.
@@ 7 Conclusion
As the first self-referential agent, Gödel Agent has to construct all task-related code autonomously, which poses significant challenges. Consequently, this work does not compare directly with the most complex existing agent systems, such as OpenDevin (Wang et al., 2024b), which have benefited from extensive manual engineering efforts. This makes it unrealistic to expect it to outperform systems that have taken researchers several months or even years to develop. The experiments presented in this paper are intended to demonstrate the feasibility of recursive self-improvement. A more robust and advanced implementation of the Gödel Agent is anticipated, with numerous potential improvements outlined in Appendix E.
@@ Appendix D Cost of Experiments
Appendix E Discussions and Future Directions
There is significant room for improvement in the effectiveness, efficiency, and robustness of the Gödel Agent’s self-improvement capabilities, which requires better initial designs. The following are some promising directions for enhancement: 1) Enhanced Optimization Modules: Utilize human priors to design more effective optimization modules, such as structuring the improvement algorithms based on reinforcement learning frameworks. 2) Expanded Modifiability: Broaden the scope of permissible modifications, allowing the agent to design and execute code that can fine-tune its own LLM modules. 3) Improved Environmental Feedback and Task Sequencing: Implement more sophisticated environmental feedback mechanisms and carefully curated task sequences during the initial optimization phase to prime the agent’s capabilities. Once the agent demonstrates sufficient competence, it can then be exposed to real-world environments.
Theoretical Analysis  Explore whether the Gödel Agent can achieve theoretical optimality and what the upper bound of its optimization might be. Determine whether the optimization process could surpass the agent’s own understanding and cognitive boundaries, and if so, at what point this might occur.
Safety Considerations  Although the current behavior of FMs remains controllable, as their capabilities grow, fully self-modifying agents will require human oversight and regulation. It may become necessary to limit the scope and extent of an agent’s self-modifications, ensuring that such modifications occur only within a fully controlled environment.
Appendix F Additional Novel Policies Designed by Gödel Agent
@@
Appendix G Comparison Between Random Sampling and Gödel Agent Performance

### v4

@@ Abstract
The rapid advancement of large language models (LLMs) has significantly enhanced the capabilities of agents across various tasks. However, existing agentic systems, whether based on fixed pipeline algorithms or pre-defined meta-learning frameworks, cannot search the whole agent design space due to the restriction of human-designed components, and thus might miss the more optimal agent design. In this paper, we introduce Gödel Agent, a self-evolving framework inspired by the Gödel machine, enabling agents to recursively improve themselves without relying on predefined routines or fixed optimization algorithms. Gödel Agent leverages LLMs to dynamically modify its own logic and behavior, guided solely by high-level objectives through prompting. Experimental results on multiple domains demonstrate that implementation of Gödel Agent can achieve continuous self-improvement, surpassing manually crafted agents in performance, efficiency, and generalizability.
@@
Xunjian Yin , Xinyi Wang , Liangming Pan , Li Lin Xiaojun Wan , William Yang Wang Peking University  University of California, Santa Barbara   University of Arizona {xjyin,wanxiaojun}@pku.edu.cnwilliam@cs.ucsb.edu
@@ 1 Introduction
There is a large body of work proposing agents capable of self-refinement. Some agents are designed to iterate over a fixed routine consisting of a list of fixed modules, while some of the modules are capable of taking self- or environment feedback to refine their actions  (Chen et al., 2023b; Qu et al., 2024a; Tang et al., 2025). This type of agent, referred to as Hand-Designed Agent, is depicted as having the lowest degree of freedom in Figure 2. More automated agents have been designed to be able to update their routines or modules in some pre-defined meta-learning routine, for example, natural language gradients (Zhou et al., 2024), meta agent (Hu et al., 2024), or creating and collecting demonstrations  (Khattab et al., 2023). This type of agent, known as Meta-Learning Optimized Agents, is depicted as having the middle degree of freedom in Figure 2. However, there are inevitably some human priors involved in these agent designs that cannot be improved during the inference time.
@@
In this paper, we choose to implement it by letting it manipulate its own runtime memory, i.e., the agent is able to retrieve its current code in the runtime memory and modify it by monkey patching (Bimal, 2012), which dynamically modifies classes or modules during execution. To allow it to update the logic of the running main function, unlike the loop-iterative approach of traditional agents, we implement the main function as a recursive function. In this function, LLM analyzes and makes a series of decisions, including reading and modifying its own code from runtime memory (self-awareness [^1] In this paper, self-awareness means that the agent can introspect and read its own code and files, not to imply any philosophical sense of consciousness or awareness. and self-modification), and interacting with the environment to gather feedback. The agent then proceeds to the subsequent recursive depth and continues to optimize itself.
To validate the effectiveness of Gödel Agent, we conduct experiments on multiple domains including coding, science, math, and reasoning. Our results demonstrate that Gödel Agent achieves significant performance gain across various tasks, surpassing various widely-used agents that require human design. The same implementation of Gödel Agent can easily adapt to different tasks by only specifying the environment description and feedback mechanism. Additionally, the case study of the optimization progress reveals that Gödel Agent can provide novel insights into agent design. Our codes are released to facilitate future research [^2] https://github.com/Arvid-pku/Godel_Agent.
@@ 4.2 Additional Designs
Table 1: Results of three paradigms of agents on different tasks. The highest value is highlighted in bold, and the second-highest value is underlined. Gödel-base is the constrained version of Gödel Agent, allowing for fair comparisons with other baselines. Gödel-free represents the standard implementation without any constraints, whose results are italicized. We report the test accuracy and the 95% bootstrap confidence interval on test sets [^4] The results of baseline models are refer to Hu et al. (2024)..
@@ 6.3 Case Study: Game of 24
The findings indicate that stronger initial policies lead to faster convergence, with smaller optimization margins, as Gödel Agent reaches its performance limit without further enhancing its optimization capabilities. Conversely, weaker initial methods result in slower convergence and larger gains, with Gödel Agent making more modifications. However, even in these cases, Gödel Agent does not outperform the results achieved using ToT. Given the current limitations of LLMs, it is challenging for Gödel Agent to innovate beyond state-of-the-art algorithms. Improvements in LLM capabilities are anticipated to unlock more innovative self-optimization strategies in the future.
7 Discussions and Future Directions
Human Self-Referential Agent
Intelligent Module brain LLM
Perceptual and Action Module body code and tool
Self-Referential Feature Humans can train their brain and body to improve, thus becoming better Self-referential agents can modify their code, even the underlying LLM, to improve themselves
Self-Awareness Question Can the brain recognize itself as a brain? Can it perceive its own mode? Can LLM understand that it is one part of the modified codes?
Table 3: An analogy of self-reference for both humans and agents
7.1 Discussions
Table 3 draws an analogy between human self-reference and the potential for self-referential capabilities in artificial agents. Inspired by this analogy, we believe that self-reference constitutes a foundational and indispensable attribute for the development of AGI, and that future agents should inherently be self-referential. As foundation models grow in power, agents can more effectively enhance their own capabilities, ultimately evolving beyond the boundaries (or limitations) of human design.
Furthermore, when an agent adjusts its own code based on feedback, this is akin to an executable version of test-time computing. In the context of LLMs, test-time computing typically involves generating additional tokens during inference, which then serve as a prefix to the final answer. This is because LLMs process information solely through text, making this their primary method for increasing computational effort at test time. For agents, however, their ability to call tools and execute code allows for far more diverse forms of test-time computing. Gödel Agent actualizes these more diverse forms of test-time computing precisely by modifying its own runtime code during test time.
7.2 Future Directions
There is significant room for improvement in the effectiveness, efficiency, and robustness of the Gödel Agent’s self-improvement capabilities, which requires better initial designs. The following are some promising directions for enhancement: 1) Enhanced Optimization Modules: Utilize human priors to design more effective optimization modules, such as genetic algorithms and reinforcement learning frameworks. 2) Expanded Modifiability: Broaden the scope of permissible modifications, allowing the agent to design and execute code that can fine-tune its own LLM modules. 3) Improved Environmental Feedback and Task Sequencing: Implement more sophisticated environmental feedback mechanisms and carefully curated task sequences during the initial optimization phase to prime the agent’s capabilities. Once the agent demonstrates sufficient competence, it can then be exposed to real-world environments.
Theoretical Analysis  Explore whether Gödel Agent can achieve theoretical optimality and what the upper bound of its optimization might be. Determine whether the optimization process could surpass the agent’s own understanding, and if so, at what point this might occur.
Safety Considerations  Although the current behavior of FMs remains controllable, as their capabilities grow, fully self-modifying agents will require human oversight and regulation. It may become necessary to limit the scope and extent of an agent’s self-modifications, ensuring that modifications occur only within a controlled environment.
@@ 8 Conclusion
As the first self-referential agent, Gödel Agent has to construct all task-related code autonomously, which poses significant challenges. Consequently, this work does not compare directly with the most complex existing agent systems, such as OpenDevin (Wang et al., 2024b), which have benefited from extensive manual engineering efforts. This makes it unrealistic to expect it to outperform systems that have taken researchers several months or even years to develop. The experiments presented in this paper are intended to demonstrate the feasibility of recursive self-improvement.
Additionally, as the agent system becomes increasingly complex through self-optimization, it may require exponentially more intelligence to understand itself. Consequently, a system capable of complete self-referential at the outset may lose this capability as it evolves (Yampolskiy, 2015). The exact point at which the agent can no longer comprehend and improve itself has not been thoroughly explored. Investigating this phenomenon, both experimentally and theoretically, could provide valuable insights into the limitations of recursive self-improvement. A more robust and advanced implementation of the Gödel Agent is anticipated, with numerous potential improvements outlined in Section 7.
Ethics Statement
Gödel Agent, capable of reading and modifying its own code, offers significant potential for advancing AI autonomy and innovation. However, this capability raises ethical and safety concerns that must be addressed to prevent harmful outcomes.
Self-modification may lead to unpredictable behavior, such as errors or unintended outputs that could violate ethical principles or produce harmful results. To mitigate these risks while preserving innovation, we propose: (1) Sandboxed Environment: Modifications should occur in an isolated sandbox to prevent unintended impacts and allow safe testing. (2) Constrained Modifications: Clear rules should limit the scope of changes to ensure safety without stifling creativity.
Further research is needed to balance safety and innovation, ensuring self-modifying agents operate within ethical boundaries. Sandboxed execution and ongoing scrutiny will help maximize benefits while minimizing risks.
@@ Appendix D Cost of Experiments
Appendix E Additional Novel Policies Designed by Gödel Agent
@@
Appendix F Comparison Between Random Sampling and Gödel Agent Performance